Problem 1. A semiconductor company has the following costs associated with Product X Fixed cost of $200M ($200,000,000) wafer yield of90% Die yield of 65% (65% of dies on a wafer make it to the next level) 12-inch wafer manufacturing cost of S5,000 per wafer Die size 12mmx12mm Testing and packaging cost of $4per die A final test yield of85% Product volume of 400M chips a) What is the cost per IC (considering both fixed and variable costs)? b) Suppose that you are a strategic planner for this company and you need to reduce the cost to make profit. Which one is a better strategy? The auswers should be quantitatively justified to get credit! Strategy 1: Reduce the fixed operating cost by 20% Strategy 2: Spend an extra Research and Development cost of $200,000,000 to improve the die yield from 65% to 80%.
Answer #1

Answer(a)

Wafer manufacturing cost as given=5000 per wafer

wafer yield=90%=5000×90%=4500

Die yield= 65%=4500×65%=2925

Testing and packaging=4 per die= 2925×4=$11700

Test yield=85%=11700*85%=9945

variable cost =9945

fixed cost=$200M

volume= 400M chips

fixed cost per chips=200M/400m=0.5M dollar= 500000 dollar

Total cost per Ic chip= fixed+ variable=500000+9945=509945.

(b)

Reduce operating cost by 20%

Fixed cost=200M×(100-20%)=200M×80%=160M

fixed cost per chips=160M/400m=0.4M dollar= 400000 dollar

Total cost per ic now= fixed+ variable=400000+9945=409945

startegy2

Extra fixed research and development expenditure= 200M

now total= 200M+200M= 400M

fixed cost per chips=400M/400m=1dollar= 1000000 dollar

Total cost=1000000+9945=1009945

yield now=1009945×80%=807956

strategy 1 should be selected.because cost is less as compared to startegy2